OpenStudy (anonymous):

what is the fifth term of the geometric sequence a1=120, a2=36, a3=10.8, a6=0.2916

OpenStudy (***[isuru]***):

can u find the common ratio ?

OpenStudy (***[isuru]***):

if the common ratio is (r) \[r = \frac{ a_{2} }{ a_{1} }\]

OpenStudy (anonymous):

0.3

OpenStudy (***[isuru]***):

so.. if u multiplied the fifth term by the common ration what will u get ?

OpenStudy (anonymous):

im trying to find the 5th

OpenStudy (***[isuru]***):

nope just thin the fifth term as a5 \[a_{5} \times r = a_{6}\] u will get the sixth term if u multiply the fifth by the common ratio got it ?

OpenStudy (***[isuru]***):

u know the value of a_6 = 0.2916 and r = 0.3 so... \[a_{5} = \frac{ a_{6} }{ r}\] \[a_{5} = \frac{ 0.2916 }{ 0.3 }\] can u complete the rest ?

OpenStudy (anonymous):

its 0.972

OpenStudy (***[isuru]***):

yep! thats the term u want !

OpenStudy (anonymous):

thank you so much. :-)

OpenStudy (***[isuru]***):

u r welcome!